Sarah has been selling 5,000 baseball caps per month at $7.50 each. When she decreased the price to $4.50 she sold 6,000baseball hats.

a.What is the demand elasticity?

b.If Sarah's marginal cost is $7 per cap, what is her desired markup and what is her initial actual markup?

c.Was decreasing the price profitable for Sarah's baseball caps business? Explain.
Answer #1

a.

demand elasticity = % change in demand/%change in price

% change in demand =( (6000 - 5000)/( (6000+5000)/2))*100 = 18.18%

% change in price = ( (4.5 - 7.5)/( (7.5+4.5)/2))*100 = -50%

demand elasticity = 18.18%/-50% = -0.3636

demand is inelastic as demand elasticity <1

b)

Initial actual markup =( (7.5-7)/7.5)*100 = 6.67%

Desired markup = - 1/price elasticity = -1/-0.3636 = 2.75%

c)

As the demand is inelastic because demand elasticity is <1, if the price is lowered, increase in quantity would not offset reduction in price thus resulting in less revenue

So decreasing price is not profitable for Sarah's baseball caps business